What is the answer to this equation 2x/5 = 4+2/3 ? If Not the answer how do i solve this equation?
Answer by mananth(16946)

2x/5 =4+2/3
the denominator has to be removed
first multiply by 3
6x/5 = 12+2
6x/5 =14
multiply by 5
6x=70
x= 70/6
x=11.66